A number is selected randomly from a container containing all the integers from 5 to 35. find P(odd | greater than 20)
1
Expert's answer
2020-11-25T19:11:16-0500

By definition,

"\displaystyle P(A|B) = \frac{P(AB)}{P(B)}"

where "P(AB)" is the probability that number is odd and greater than 20, "P(B)" - number is greater than 20.

There are 31 integer number from 5 to 35. Among them, there are 15 numbers that are greater than 20. The amount of numbers that are odd and greater than 20 is 7. Therefore,

"\displaystyle P(AB) = \frac{7}{31}, \; P(B) = \frac{15}{31}"

and

"\displaystyle P(A|B) = \frac{P(AB)}{P(B)}=\frac{7}{15}"

Answer: 7/15